OpenStudy (bbb911):

What are the ingredients for one-half of the recipe? 6 1/2 cups of flour, 7 teaspoons of cinnamon, 1/2 teaspoon of salt. What are the ingredients for one-half of the recipe?

OpenStudy (anonymous):

take these mixed numbers first and switch them into fractions to do this take your whole number and multiply by the denominator of the fraction. In simplicity, You have 6 cups, and it takes 2 one-half cups to make one full cup, so to find out how many half cups 6 cups are, you just doo \[6*2=12=\frac{12}{2}=12(\frac{1}{2})\]

OpenStudy (anonymous):

let me rephrase that \[6*\frac{2}{2}=\frac{12}{2}=12(\frac{1}{2})\]

OpenStudy (anonymous):

so in other words 6 cups will make you 12 on half cups

OpenStudy (bbb911):

its says i should multiply 6 1/2 and 1/2 right

OpenStudy (anonymous):

not exactly you want to do this step by step 1) Take the whole number . In the first case you have 6 cups 2) Multiply this number by 1 (using the denominator of the preferred units: in this case one half cups. \[1=\frac{2}{2}\]*This has one half on the bottom. This is what we need!. but if we divide by two we need to multiply by two otherwise we'll screw it up =/ 3)This should get you your whole number in terms of onehalf cups

OpenStudy (anonymous):

does this make sense?

OpenStudy (bbb911):

well actallu 6 and 1/6 cups

OpenStudy (anonymous):

alright so what has the denominator of 6 but is still one

OpenStudy (anonymous):

aka \[1=\frac{?}{6}\]

OpenStudy (bbb911):

ohok so 6/1 right

OpenStudy (anonymous):

does \[1=\frac{6}{1}\]

OpenStudy (anonymous):

oh wait do you mean the numerator has to be 6?

OpenStudy (anonymous):

did you mean this \[1=\frac{6}{6}\]

OpenStudy (bbb911):

cos im a little confused on like how it is like 6 1/6 then 7 then 1/2 cos the 1/6 is the fraction and the whole number is 6 which is included with the 1/6 fraction . hopefully that made sense

OpenStudy (anonymous):

well the others are pretty simple ... so lets do those first

OpenStudy (bbb911):

kk

OpenStudy (anonymous):

So we have 7 teaspoons of cinnamon, and we only need half of that. If we multiply by one-half (or divide by 2) \[\frac{1}{2}*7=\frac{7}{2}\] This fraction can be simplified into a mixed number because 2 goes into seven 3 times with a remainder of 1/2 \[3+\frac{1}{2}\]

OpenStudy (anonymous):

so you'll need three and one-half teaspoons of cinnamon

OpenStudy (bbb911):

then for the flour?

OpenStudy (anonymous):

next you need half of a half teaspoon of salt. using the same technique before, divide by 2 of multiply by 1/2. You'll get 1/4 teaspoon ... this can also be shown as this 

OpenStudy (anonymous):